Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month's budget appear below:



Selling price $27 per unit
Variable expenses $16 per unit
Fixed expenses $8,910 per month
Unit sales 960 units per month


Requirement 1:
Compute the company's margin of safety. (Omit the "$" sign in your response.)

Margin of safety $


Requirement 2:
Compute the company's margin of safety as a percentage of its sales. (Round your answer to the nearest whole percent. Omit the "%" sign in your response.)


Margin of safety as a percentage of sales %
Answer #1
Concepts and reason

CVP analysis: CVP is one of the techniques of decision accounting to achieve the targeted profits by changing different variables. The relationship between cost, profit and sales volume provides the basis for the manager to take effective steps about the future profits.

Sales revenue: The amount earned by selling goods or services is known as sales revenue. This is the main source of revenue for a company. This is usually referred to as gross sales revenue. Gross sales revenue less returns, allowances, and discounts are known as net sales.

Fundamentals

Contribution margin: The difference between the sales revenue and the variable costs is called contribution margin. The management uses contribution margin to develop the weight of sales mix for multiple products. The contribution margin signifies the profit earned after deducting the fixed costs.

Selling price: Selling price is the price at which products sells in the market.

Breakeven point: The point of sales at which there is no profit or loss is called as Break Even Point (BEP). It means the total cost equals the total revenue at this point.

Variable costs: These costs vary with the number of units produced or for the services provided. For example, the labor costs increase if the number of labor hours is increased, and the labor costs decrease if the number of labor hours is decreased.

Fixed cost: These are the costs which remain constant throughout the process of manufacturing or for the services rendered. They are incurred irrespective of number of units produced.

Margin of safety: Margin of safety is an excess of budgeted sales than break-even sales.

Margin of safety ratio: Margin of safety ratio is calculated by dividing margin of safety by expected sales. It is also called as margin of safety as a percentage of sales.

1)

Calculate margin of safety.

Marginofsafety=ExpectedsalesBreakevensales=$25,920$21,870=$4,050\begin{array}{c}\\{\rm{Margin of safety = Expected sales }} - \;{\rm{Break even sales}}\\\\{\rm{ = \$ 25,920 -- \$ 21,870}}\\\\{\rm{ = \$ 4,050}}\\\end{array}

Therefore, the margin of safety is $4,050.

Working note:

Calculate expected sales.

Expectedsales=Unitsales×Sellingpriceperunit=960×$27=$25,920\begin{array}{c}\\{\rm{Expected sales = Unit sales }} \times {\rm{ Selling price per unit }}\\\\{\rm{ = 960 }} \times {\rm{ \$ 27 }}\\\\{\rm{ = \$ 25,920}}\\\end{array}

Calculate break-even sales.

Breakevensales=((Fixedexpenses(SellingpriceperunitVariableexpensesperunit))×Sellingpriceperunit)=(($8,910($27$16))×$27)=$21,870\begin{array}{c}\\{\rm{Break even sales = }}\left( \begin{array}{l}\\\left( {\frac{{{\rm{Fixed expenses}}}}{{\left( \begin{array}{l}\\{\rm{Selling price per unit }}\\\\ - {\rm{ Variable expenses per unit}}\\\end{array} \right)}}} \right){\rm{ }}\\\\ \times {\rm{ Selling price per unit}}\\\end{array} \right)\\\\{\rm{ = }}\left( \begin{array}{l}\\\left( {\frac{{{\rm{\$ 8,910}}}}{{\left( {{\rm{\$ 27}}\, - \;{\rm{\$ 16}}} \right)}}} \right){\rm{ }}\\\\ \times {\rm{ \$ 27}}\\\end{array} \right)\\\\ = \;\$ 21,870\\\end{array}

2)

Calculate margin of safety as a percentage of sales.

Marginofsafetyasapercentageofsales}=(MarginofsafetyExpectedsales)×100=($4,050$25,920)×100=15.63%\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Margin of safety as a }}\\\\{\rm{percentage of sales}}\\\end{array} \right\}{\rm{ = }}\left( {\frac{{{\rm{Margin of safety}}}}{{{\rm{Expected sales}}}}} \right){\rm{ }} \times {\rm{ 100 }}\\\\{\rm{ = }}\left( {\frac{{{\rm{\$ 4,050}}}}{{{\rm{\$ 25,920}}}}} \right){\rm{ }} \times {\rm{ 100 }}\\\\{\rm{ = 15}}{\rm{.63\% }}\\\end{array}

Therefore, the margin of safety as a percentage of sales is 15.63%.

Ans: Part 1

Margin of safety is $4,050.

Part 2

Margin of safety as a percentage of sales is 15.63%.